A 6-cylinder car is driven 37,500 miles in 3 years. At that rate, in how many years will the car have 100,000 miles on it?
Answer by checkley75(3666)

37,500/3=100,000/X CROSS MULTIPLY:
17,500X=300,000
X=300,000/17,500
X=17.143 YEARS